OpenStudy (anonymous):

How do you apply the power rule when E is raised by an power to a power? f[x] = e^(x^2) f'[x] = ???

TheSmartOne (thesmartone):

@Michele_Laino QH help :)

OpenStudy (amistre64):

derivative or integration?

OpenStudy (mathstudent55):

You don;t use the power rule here. Here you need the derivative of e^u, where u is a function in x.

OpenStudy (amistre64):

i see, derivative, chain and then power

OpenStudy (anonymous):

trying to get to derivative.. yes

OpenStudy (mathstudent55):

\(\dfrac{d}{dx} e^u = e^u \dfrac{d}{dx}u\)

OpenStudy (michele_laino):

here we have to apply the derivation of composed functions

OpenStudy (mathstudent55):

Oh, I see. The power rule does come in to differentiate the exponent, x^2.

OpenStudy (anonymous):

so its like 2 functions?

OpenStudy (michele_laino):

noe we can write your function as follows: \[f\left( x \right) = {e^{g\left( x \right)}},\quad g\left( x \right) = {x^2}\]

OpenStudy (michele_laino):

so the requested first derivative is given by the subsequent formula: \[f'\left( x \right) = \frac{{df}}{{dg}} \times \frac{{dg}}{{dx}} = {e^{g\left( x \right)}} \times 2x = 2x{e^{{x^2}}}\]

OpenStudy (anonymous):

wicked .. thank you now I just need to get my head around that.. :) ... first day solving derivative problems using rules. .. :)

OpenStudy (michele_laino):

thank you! :)

OpenStudy (anonymous):

makes sense though :)

OpenStudy (michele_laino):

need more help?

OpenStudy (anonymous):

all good thnx

OpenStudy (michele_laino):

thanks!! :)